Mitch is at the top of Grizzly ski run. His brother bill is looking up the hill from the ski lodge which 200 feet away at an angle of 45. What is the height of the ski run?

Answer by checkley75(3666)

THE HYPOTENUSE OF THIS TRIANGLE IS 200 FEET THUS THE EQUATION IS
X^2+X^2=200^2
2X^2=200^2
2X^2=40000
X^2=40000/2
X^2=20000
X=141.42 FEET IS THE HEIGHT OF THE HILL.